OpenStudy (anonymous):

Solve the system 2x - 3y = 0 and x + 4y = -11 by graphing each line

OpenStudy (anonymous):

Graph both of them wherever they intersect the x and y coordinates are your answer :D

OpenStudy (anonymous):

numerically if you want to check your answer you can use substitution method. so 2x = 3y and x = 3y/2 so then 3y/2 + 4y = -11 so 3y/2 + 8y/2 = 11y/2 = -11 so 11y = -22 and y = -2 so then x = 3(-2)/2 = -3

OpenStudy (anonymous):

so (-3,-2) is where the two lines would intersect